Deduce the final structure consistent to the 1H-NMR data.

C5H9BrO2

1.2 ppm (triplet, 3H), 2.9 ppm (triplet, 2H), 3.5 ppm (triplet, 2H), 4.0 ppm (quartet, 2H)
Answer #1

In DU, p is the no of carbon atoms and n is the no of Hydrogen atoms.

Home a-12, triplet, 3H b-4.0 quartet, 2H c-2.9, triplet, 2 2. B o Cg Hq BYO DU = 2(5)+2-9+) - 12-2-1 DU = 2(8) 72-(n+1) tripl